OpenStudy (anonymous):

A stone that starts at rest is in free fall for 8.0 s. a. Calculate the stone’s velocity after 8.0 s b. What is the stone’s displacement during this time?

OpenStudy (anonymous):

This is my way of doing this a. 80m/s b. d = 1/2at^2 = 5*64 = 320m